OpenStudy (anonymous):

The index of refraction of a substance is 1.5. Find the sine of the angle of refraction if the sine of the angle of incidence is 0.15. sin r = _____

OpenStudy (anonymous):

@ShailKumar

OpenStudy (anonymous):

Do you know about Snell's law ?

OpenStudy (anonymous):

No :l

OpenStudy (anonymous):

OK. Snell's law is \[\frac{sin _\ i}{sin _\ r} = refractive\ index\]

OpenStudy (anonymous):

 i = angle of incidence and r = angle of refraction

OpenStudy (anonymous):

Can you find sin r now ?

OpenStudy (anonymous):

No and no :/ 1 im still confused on how its done and 2 I dont have a calculator

OpenStudy (anonymous):

I had one yesterday and i tried working it out and i had gotten 10

OpenStudy (anonymous):

Using Snell's law sin r = sin i / refractive index So, sin r = 0.15/1.5 = 1/10 = 0.1 is the answer

OpenStudy (anonymous):

Okay it makes a bit more sense and im sorry for not being anything prepared at all :/

OpenStudy (anonymous):

It's ok :)

OpenStudy (anonymous):

Thank you :D